The number of hours of sleep each student gets per night is approximately normally distributed with mean of 7 hours and standard deviation of 1.25 hours.

(c) Cognitive functioning drops off dramatically without enough sleep. Suppose that Health Services wants to help the students who sleep the least. Due to budget constraints, they will focus their efforts on students who are in the bottom 5% of hours slept. At what number of hours should they focus their efforts?
Answer #1

Question is answered below.

u=7, J=1.25 z = X-y - X-7 1-25 ) - 0.05oo. 1.25 X-7 u -1.65 1.25 x = 7- 1.65X 1.25 4.9375 45. X - 5 hours approximately